OpenStudy (anonymous):

Sharon went to solve the following system and her steps are shown. x - y = 9 x + y = 15 Step One: 2x = 26 Add the systems Step Two: x = 13 Solve for x Step Three: 13 + y = 15 y = 2 Solve for y Was she correct? Justify your answer. If incorrect, describe the error.

OpenStudy (dan815):

Plug the x and y values into your system of equations to confirm

OpenStudy (campbell_st):

well what happens if you add the 2 equations x - y = 9 x + y = 15 ========== 2x = ?

OpenStudy (dbzfan836):

hmmm, I think x=3

OpenStudy (dbzfan836):

( ͡° ͜ʖ ͡°)

OpenStudy (anonymous):

really? @Dbzfan836

OpenStudy (triciaal):

no, her method is correct but she made a calculation error 9 + 15 = 24 and not 26